# Sequence Generation Modeling for Continuous Value Prediction

### v2

@@ Abstract.
Continuous value prediction (CVP) plays a crucial role in short video recommendation, capturing user preferences through precise numerical estimations. However, traditional regression-based methods often struggle with challenges like wide value ranges and imbalanced data, leading to prediction bias. While ordinal classification approaches have been introduced to address these issues, their reliance on discretization reduces accuracy and overlooks inherent relationships between intervals. To overcome these limitations, we introduce a novel Generative Regression (GR) framework for CVP, inspired by sequence generation techniques in language modeling. Our method transforms numerical values into token sequences through structural discretization, preserving original data fidelity while improving prediction precision. Leveraging a carefully crafted vocabulary and label encoding, GR employs curriculum learning with an embedding mixup strategy to bridge training-inference gaps. Experimental evaluations on four public datasets and one large-scale industrial dataset validate the superiority of GR over existing methods. Real-world A/B tests on Kuaishou, a leading video platform, further demonstrate its practical effectiveness. Additionally, GR proves adaptable to other regression tasks, such as Lifetime Value (LTV) prediction, showcasing its potential as a robust solution for diverse CVP challenges.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06submissionid: 6719ccs: Information systems Recommender systems
@@ 1. Introduction
In contrast to limited and discrete actions such as liking, following, and sharing, watch time generally exhibits a wide range and long-tailed distribution. This characteristic renders it fundamentally a regression problem for prediction. However, the continuous nature of watch time also makes it vulnerable to outliers and potential predictive bias, thereby increasing modeling complexity and challenges of accurate prediction. Some methods (Zhan et al., 2022; Zhao et al., 2024; Zheng et al., 2022; Zhao et al., 2023a) optimize watch time prediction from a debiasing perspective but have not yet adequately addressed the core challenges of regression. Some others (Sun et al., 2024; Lin et al., 2023) transform the prediction problem into an Ordinal Regression (OR) task by employing a series of binary classifications across various predefined time intervals, as separately shown in Fig. 1(a) and Fig. 1(b). Nevertheless, such a modeling paradigm still exhibits two major limitations as follows:
Firstly, conditional dependencies among time intervals are not fully leveraged. The dependencies are solely reflected in the definition of the labels, predictions across different time intervals are often produced independently, thereby hindering the potential for effective error correction.
Secondly, the ordinal regression paradigm decomposes the regression target into a series of binary classification problems defined by fixed time intervals, which inherently confines the labeling of samples within these rigid divisions. As a result, model performance is highly contingent on the method of time interval segmentation. Coarse segmentation, i.e., wider spans, may overlook subtle differences, while overly fine segmentation, i.e., narrower spans, can result in insufficient data, degrading the model’s ability to learn effective features (Sun et al., 2024).
@@
On the other hand, unlike ordinal regression methods (Sun et al., 2024; Lin et al., 2023) that restrict outputs to binary classification within fixed time intervals, our GR model provides the flexibility of allowing each predictive step to choose from a vocabulary of tokens, where each token represents a distinct time slot in positive real number space. This enlarged solution space enables GR to generate a broader set of potential sequences, thereby improving its capacity to generalize across diverse watching behaviors and leading to more accurate and personalized predictions. A comprehensive comparison of GR with other regression methods is shown in Tab. 1.
For token definition and watch time segmentation, we propose a data-driven unified vocabulary construction method, which mitigates token imbalance and eliminates reliance on manual design, and a label encoding strategy allows a lossless restoration of watch time values, thereby enhancing the model’s generality and generalization capability. To accelerate model convergence, we adopt curriculum learning (Bengio et al., 2009) strategy that combines teacher forcing (Venkatraman et al., 2015) with standard training to alleviate training-and-inference inconsistency, commonly known as exposure bias (Venkatraman et al., 2015; Ding and Soricut, 2017). Besides, leveraging our insights into the training process, we propose an embedding mixup method to compensate for output-to-input gradients. This approach enhances model performance at a lower computational cost by leveraging the semantic additivity of tokens while ensuring consistency between training and inference.
Table 1. A comprehensive comparison with existing deep regression methods. ‘H’, ‘M’, and ‘L’ indicate high, medium, and low respectively. Complexity represents computation overhead during inference, with $m$, $n$, $d$, and $t$ indicating the last hidden feature size, discretization segments, vocabulary size, and sequence length, respectively.
Methods Predictive Paradigm Discre -tization Accuracy Versatility Complexity
VR Discriminative - L H $\mathcal{O}(m)$
TPM (Lin et al., 2023) Discriminative Tree-based M M $\mathcal{O}(mn)$
CREAD (Sun et al., 2024) Discriminative Linear M M $\mathcal{O}(mn)$
GR Generative Sequential H H $\mathcal{O}(mdt)$
@@
    To enhance the generality and adaptability of GR, we develop a data-driven unified vocabulary design and a label encoding method. Additionally, we introduce curriculum learning with embedding mixup to mitigate exposure bias and accelerate model training.
@@
    Extensive online and offline experiments demonstrate that GR significantly outperforms existing SOTA models. We further analyze the impact of key factors like vocabulary design and model architecture, to provide a clear understanding of the mechanisms and processes underlying GR.
@@ 2. Related Work
Watch time prediction aims to estimate the video watch time based on the user’s profile, historical interactions, and video characteristics. Value regression (VR) directly predicts the absolute value of watch time, assessing model accuracy by mean square error (MSE). WLR (Covington et al., 2016) applies watch duration as weights within its loss function, yet it may suffer from significant bias due to its weighting mechanism. D2Q (Zhan et al., 2022) segments data by duration, predicting watch time quantiles within these segments to alleviate duration bias. CWM (Zhao et al., 2024) introduces a counterfactual watch time, estimating a video’s hypothetical full watch time to gauge user interest. D2Co (Zhao et al., 2023a) differentiates actual user interest from duration bias and noisy watching using a duration-wise Gaussian mixture model. However, these methods have not yet adequately addressed the core challenges of regression. Some solutions transform the problem into ordinal regression. CREAD (Sun et al., 2024) introduces an error-adaptive discretization technique to construct dynamic time intervals. TPM (Lin et al., 2023) utilizes hierarchical labels to model relationships across varying granularities of time intervals. Yet, these approaches are unable to fully capitalize on the interdependencies among these intervals and heavily rely on time interval segmentation.
@@
Ordinal regression, also known as ordinal classification, is a type of predictive modeling strategy employed when the outcome variable is ordinal and the relative order of labels is important, such as age prediction (Niu et al., 2016), monocular depth perception (Fu et al., 2018), and head-pose estimation (Hsu et al., 2018). CNNOR (Liu et al., 2017) introduces a deep learning model for ordinal regression tailored to small datasets. SORD (Diaz and Marathe, 2019) employs a soft ordinal label approach to train the ordinal regression model. Additionally, (Li et al., 2021) crafts a probabilistic embedding strategy specifically for ordinal regression tasks. Ordinal regression has not been applied to watch time prediction until the introduction of CREAD (Sun et al., 2024) and TPM (Lin et al., 2023). These models follow the traditional approach of decomposing regression into multiple binary classification tasks, achieving significant benefits.
@@
Figure 2. The framework of GR model, which adopts an encoder-decoder RNN-based architecture. The encoder extracts user and video features, while the decoder predicts watch time in an autoregressive manner and employs the curriculum learning with embedding mixup (CLME) strategy to alleviate training-and-inference inconsistency introduced by teacher forcing.
@@
Sequence generation requires the model to understand the context of an input sequence and generate an appropriate output sequence. Early researches focus on natural language processing (NLP) tasks such as machine translation (Sutskever, 2014; Elman, 1990; Graves and Graves, 2012; Cho, 2014), text summarization (Bahdanau, 2014; Vaswani, 2017), and dialogue generation (Devlin, 2018; Raffel et al., 2020). In recommendation systems, sequential recommendation methods have been proposed to capture sequential patterns. GRU4Rec (Hidasi, 2015) is a session-based recommendation model with GRU. SASRec (Kang and McAuley, 2018) utilizes a self-attention mechanism to capture both long-term and short-term user preferences. BERT4Rec (Sun et al., 2019) employs a bidirectional transformer to encode item sequences. However, sequence generation models in recommendation systems predominantly focus on predicting the sequence of user behaviors, and their application to watch time prediction remains unexplored.
@@ 3.1. Problem Formulation
Given a dataset $\mathcal{D}=\left\{(\bm{u_{i}},\bm{v_{i}},y_{i})\right\}_{i=1}^{N}$, where $\bm{u_{i}}$ and $\bm{v_{i}}$ represent the user-side features (such as user ID, static profile, and historical behaviors etc.) and the item-side (videos in this paper) features (e.g. tags, duration and category) of the $i$-th example respectively  [^1] We omit the context-side features for simplicity., $y_{i}\in\mathbb{R}$ is the corresponding watch time of the $i$-th example collected from recommendation system logs. Value regression (VR) methods aim to learn a function $f(\cdot)$ that directly maps the input features to a real-valued output, i.e., $y_{i}=f([\bm{u_{i}};\bm{v_{i}}])$.
Sequence generation in GR is mostly based on autoregressive language modeling. Specifically, we introduce a vocabulary $\mathcal{V}=\left\{w_{j}\right\}_{j=1}^{V}$, where each element $w_{j}$ represents a predefined time slot (e.g. 5 seconds, 10 seconds, etc.). The details of vocabulary construction are presented in Sec. 3.3. Here, these time slots are analogous to tokens in language models (LMs). Thus, “token” and “time slot” will be used interchangeably in the sequel. The vocabulary embedding matrix is denoted as $\bm{E}\in\mathbb{R}^{V\times D}$, where $V$ is the vocabulary size, and $D$ is the dimension of the time slot embeddings.
We decompose $y_{i}$ into a sequence of tokens $\bm{s_{i}}=\{s_{i}^{1},...,s_{i}^{T_{i}}\}$, where $s_{i}^{t}\in\mathcal{V}$ and $T_{i}$ is the length of the sequence. We call this process label encoding, which is detailed in Sec. 3.4. On the other hand, we design a label decoding function $\text{g}(\cdot)$ such that derives the original watch time $y_{i}$ from $\bm{s_{i}}$, i.e., $y_{i}=\text{g}(\bm{s_{i}})=\sum_{t=1}^{T_{i}}\text{g}(s_{i}^{t})\in\mathbb{R}$. Our goal is to train a sequence generation model, given the user and video features $(\bm{u_{i}},\bm{v_{i}})$, generates the corresponding sequence of watch time slots $\hat{s}_{i}=\{\hat{s}_{i}^{1},\hat{s}_{i}^{2},...,\hat{s}_{i}^{T_{i}}\}$, and in turn, from which the predicted watch time $\hat{y_{i}}=\sum_{t=1}^{T_{i}}\text{g}(\hat{s}_{i}^{t})$ approximates the actual watch time $y_{i}$.
@@ 3.2. The Generative Regression (GR) Model
As shown in Fig. 2, GR adopts an encoder-decoder RNN-based architecture. The encoder extracts user and video features, while the decoder predicts watch time in an autoregressive manner.
@@ 3.2.1. Encoder
Unlike traditional sequence-to-sequence tasks or user behavior modeling in recommendation systems, watch time prediction does not inherently depend on history action sequences. Consequently, directly using an RNN as the encoder — designed for processing temporally dependent sequential data — is inappropriate and may result in suboptimal feature representation.
To ensure the generality and simplicity of the model, we use a feed-forward network (FFN) as the encoder. However, it can be replaced with any complex model architecture. Formally, the encoder extracts user and video features to produce a fixed-length hidden feature $\bm{h_{i}}\in\mathbb{R}^{1\times D}$ that will be fed to the decoder as follows:
@@ 3.2.2. Decoder
where $\theta$ is the model parameter and $\bm{\hat{s}_{i}^{<t}}$ is the tokens generated before. Utilizing the chain rule, the overall probability of generating the sequence can be expressed as
(3) $P_{\theta}(\bm{s_{i}}\mid\bm{h_{i}})=P_{\theta}(s_{i}^{1},...,s_{i}^{T_{i}}\mid\bm{h_{i}})=\prod_{t=1}^{T}P(s_{i}^{t}\mid\bm{h_{i}},\bm{{s}_{i}^{<t}})$
There remain three issues to be addressed: how to construct an effective vocabulary? how to encode $y_{i}$ into a sequence $\bm{s_{i}}$? and how to optimize the model? which are detailed in the following sections.
@@ 3.3. Vocabulary Construction
  • Completeness: The vocabulary $\mathcal{V}$ must be able to represent all watch time values $\{y_{i}\}_{i=1}^{N}$ using a finite number of tokens almost without loss [^2] The label space is continuous, it is impossible to recover the original watch time from a discretized sequence with zero error.. Also, each token must be unique.
@@
One intuitive strategy is to select watch time values from the dataset as tokens based on several fixed percentiles, yet failing to meet the completeness principle. An alternative is to select watch time values as tokens based on one fixed percentile, then subtract the token values from all watch time values that exceed them, repeating this process until the residuals become negligible, which fails to meet the balance principle.
@@
1:Dataset labels $\bm{Y}=\{y_{j}\}_{j=1}^{N}$, initially empty Vocabulary $\mathcal{V}=\{\}$, precision threshold $\epsilon$, start percentile $q_{\text{start}}$, end percentile $q_{\text{end}}$, decay rate $\alpha$, minimal restoration error $\epsilon$.
2:Sort $\bm{Y}$ in descending order to obtain $\hat{\bm{Y}}$.
@@
5:     Compute the $q$-percentile $z_{i}$ of $\bm{Y}$.
6:     if $z_{i}=0$ then $\triangleright$ Terminate if the percentile value is zero
@@
9:     Insert $z_{i}$ into vocabulary $\mathcal{V}$.
@@
$\hat{y}_{j}=\begin{cases}\hat{y}_{j},&\text{if }\hat{y}_{j}<z_{i},\\ \hat{y}_{j}-z_{i},&\text{otherwise}\end{cases}$
@@
To meet these principles above, we develop a data-drived vocabulary construction algorithm based on dynamic percentile, detailed in Algorithm. 1. Specifically, the algorithm starts with an initially large percentile $q_{start}$ and decreases it according to the decay rate $\alpha$, ensuring that the percentile does not fall below the endpoint $q_{end}$. This strategy expedites the reduction of tail values, rapidly decreasing the variance among updated values, which effectively mitigates the challenges posed by the long-tailed distribution in the dataset, for which we provide detailed validation in the experimental section.
Under the generative regression framework, it is important to note that the process of constructing a vocabulary and converting the task target (e.g. watch time) into a sequential token format (a.k.a label encoding) resembles the syntax construction of a language. However, we do not contend that the aforementioned strategy is the most optimal one.

### v3

# Generative Regression Based Watch Time Prediction for Short-Video Recommendation

@@ Abstract.
Watch time prediction (WTP) has emerged as a pivotal task in short video recommendation systems, designed to quantify user engagement through continuous interaction modeling. Predicting users’ watch times on videos often encounters fundamental challenges, including wide value ranges and imbalanced data distributions, which can lead to significant estimation bias when directly applying regression techniques. Recent studies have attempted to address these issues by converting the continuous watch time estimation into an ordinal regression task. While these methods demonstrate partial effectiveness, they exhibit notable limitations: (1) the discretization process frequently relies on bucket partitioning, inherently reducing prediction flexibility and accuracy and (2) the interdependencies among different partition intervals remain underutilized, missing opportunities for effective error correction.
Inspired by language modeling paradigms, we propose a novel Generative Regression (GR) framework that reformulates WTP as a sequence generation task. Our approach employs structural discretization to enable nearly lossless value reconstruction while maintaining prediction fidelity. Through carefully designed vocabulary construction and label encoding schemes, each watch time is bijectively mapped to a token sequence. To mitigate the training-inference discrepancy caused by teacher-forcing, we introduce a curriculum learning with embedding mixup strategy that gradually transitions from guided to free-generation modes.
We evaluate our method against state-of-the-art approaches on two public datasets and one industrial dataset. We also perform online A/B testing on the Kuaishou App to confirm the real-world effectiveness. The results conclusively show that GR outperforms existing techniques significantly. Furthermore, we successfully apply GR to Lifetime Value (LTV) prediction, achieving 17.66% MAE improvement over existing methods. These results validate GR as a generalizable solution for continuous value prediction tasks in recommendation systems.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
@@ 1. Introduction
In contrast to limited and discrete actions such as liking, following, and sharing, watch time generally exhibits a wide range and long-tailed distribution, making it fundamentally a regression problem for prediction. Some methods (Zhan et al., 2022; Zhao et al., 2024; Zheng et al., 2022; Zhao et al., 2023a; Zhang et al., 2023; Tang et al., 2023) optimize watch time prediction from a debiasing perspective but have not yet adequately addressed the core challenges of regression. Some others (Sun et al., 2024; Lin et al., 2023) transform the prediction problem into an Ordinal Regression (OR) task by employing a series of binary classifications across various predefined time intervals (buckets), as separately shown in Fig. 1(a) and Fig. 1(b). While effective, such a modeling paradigm still exhibits two major limitations as follows:
Firstly, conditional dependencies among time intervals are not fully leveraged, which are solely reflected in the definition of the labels. Predictions across different time intervals are often produced independently, thereby hindering the potential for effective error correction and leading to suboptimal results. We provide rigorous theoretical proof of this limitation in the supplementary material.
Secondly, the strict discretization process within fixed time intervals in ordinal regression makes model performance highly contingent on the method of time interval segmentation, inherently reducing prediction flexibility and accuracy. This approach performs binary classification across all predefined buckets, with the final prediction derived as the sum of bucket sigmoid probabilities multiplied by their corresponding bucket span values. Due to the wide range of actual watch times (Sun et al., 2024; Tang et al., 2023), tail buckets often have excessively large span values, which can disproportionately amplify prediction errors for samples with shorter watch times, even when the binary probabilities of these tail buckets are minimal. Additionally, the scrolling mode of short video platforms results in a high percentage of videos with relatively short watch times in real-world scenarios, further exacerbating the overall fitting error.
@@
On the other hand, unlike ordinal regression methods (Sun et al., 2024; Lin et al., 2023) that restrict outputs to binary classification within fixed time intervals, our GR model offers the flexibility for each predictive step to not only select from a vocabulary of tokens—each representing a distinct time slot in positive real number space, but also output an end-of-sequence (¡EOS¿) token. This flexibility enables GR to generate a broader set of potential sequences, thereby improving its capacity to generalize across diverse watching behaviors and leading to more accurate and personalized predictions.
For token definition and watch time segmentation, we propose a data-driven unified vocabulary construction method, which mitigates token imbalance and eliminates manual design reliance, and a label encoding strategy allows a lossless restoration of watch time values, thereby enhancing the model’s generality and generalization capability. To accelerate model convergence, we adopt curriculum learning (Bengio et al., 2009) strategy during training to alleviate training-and-inference inconsistency, commonly known as exposure bias (Venkatraman et al., 2015; Ding and Soricut, 2017). Besides, leveraging our insights into the training process, we propose an embedding mixup method to compensate for output-to-input gradients. This approach enhances model performance at a lower computational cost by leveraging the semantic additivity of tokens while ensuring consistency between training and inference.
@@
    To enhance generality and adaptability, we develop a data-driven unified vocabulary design and a label encoding method. Additionally, we introduce curriculum learning with embedding mixup to mitigate exposure bias and compensate for output-to-input gradients to accelerate model training.
@@
    Extensive online and offline experiments show that GR significantly outperforms existing SOTA models. We further analyze the underlying reasons for performance gain and the impact of key factors like vocabulary design to provide a clear understanding of the mechanisms underlying GR.
@@ 2.1. Watch Time Prediction (WTP)
WTP aims to estimate the video watch time based on the user’s profile, historical interactions, and video characteristics. Value regression (VR) directly predicts the absolute value of watch time, assessing model accuracy by mean square error (MSE). Subsequent WTP methods can be roughly divided into two groups. The first focuses on optimizing WTP from a debiasing perspective (Zhan et al., 2022; Zhao et al., 2024; Zheng et al., 2022; Zhao et al., 2023a; Zhang et al., 2023; Tang et al., 2023). CWM (Zhao et al., 2024) introduces a counterfactual watch time, estimating a video’s hypothetical full watch time to gauge user interest. D2Co (Zhao et al., 2023a) differentiates actual user interest from duration bias and noisy watching using a duration-wise Gaussian mixture model. However, these methods have not yet adequately addressed the core challenges of regression. The second transforms the regression task into classification (Sun et al., 2024; Lin et al., 2023; Covington et al., 2016). CREAD (Sun et al., 2024) introduces an error-adaptive discretization technique to construct dynamic time intervals. TPM (Lin et al., 2023) utilizes hierarchical labels to model relationships across varying granularities of time intervals. Yet, these approaches are unable to fully capitalize on the interdependencies among these intervals and heavily rely on time interval segmentation.
2.2. Ordinal Regression
@@
OR is a type of predictive modeling strategy employed when the outcome variable is ordinal and the relative order of labels is important, such as age prediction (Niu et al., 2016), monocular depth perception (Fu et al., 2018), and head-pose estimation (Hsu et al., 2018). Recent works include specialized architectures like CNNOR (Liu et al., 2017), alternative training paradigms using soft labels such as SORD (Diaz and Marathe, 2019), and dedicated probabilistic embedding methods (Li et al., 2021). It has not been applied to watch time prediction until the introduction of CREAD (Sun et al., 2024) and TPM (Lin et al., 2023). These models decompose the regression task into multiple binary classification tasks, achieving significant benefits.
@@
Figure 2. The framework of the GR model, which adopts an encoder-decoder architecture. The encoder extracts user and video features, while the decoder predicts watch time in an autoregressive manner and employs the curriculum learning with embedding mixup (CLEM) strategy to alleviate training-and-inference inconsistency introduced by teacher forcing.
2.3. Sequence Generation
@@
Sequence generation learns contextual sequence mappings, initially prominent in NLP for tasks like machine translation (Sutskever, 2014; Cho, 2014) and text summarization (Bahdanau, 2014; Vaswani et al., 2017). This paradigm extended to recommendation systems for capturing sequential user behavior patterns. In recommendation systems, sequential recommendation methods have been proposed to capture sequential patterns. GRU4Rec (Hidasi, 2015) is a session-based recommendation model with GRU. SASRec (Kang and McAuley, 2018) utilizes a self-attention mechanism to capture both long-term and short-term user preferences. BERT4Rec (Sun et al., 2019) employs a bidirectional transformer to encode item sequences. However, these sequential recommendation methods have predominantly focused on predicting the sequence of user behaviors, and their application to watch time prediction remains unexplored.
@@ 3.1. Problem Formulation
Given a dataset $\mathcal{D}=\left\{(\bm{u_{i}},\bm{v_{i}},y_{i})\right\}_{i=1}^{N}$, where $\bm{u_{i}}$ and $\bm{v_{i}}$ represent the user-side features (such as user ID, static profile, and historical behaviors etc.) and the item-side (videos in this paper) features (e.g. tags, duration and category) of the $i$-th example respectively   [^1] We omit the context-side features for simplicity., $y_{i}\in\mathbb{R}$ is the corresponding watch time of the $i$-th example collected from recommendation system logs. Value regression methods aim to learn a function $f(\cdot)$ that directly maps the input features to a real-valued output, i.e., $y_{i}=f([\bm{u_{i}};\bm{v_{i}}])$.
Sequence generation in GR is mostly based on autoregressive language modeling. Specifically, we introduce a vocabulary $\mathcal{V}=\left\{w_{j}\right\}_{j=1}^{V}$, where $V$ is the vocabulary size and each element $w_{j}$ represents a predefined time slot (e.g. 5 seconds, 10 seconds, etc.). The details of vocabulary construction are presented in Sec. 3.3. Here, these time slots are analogous to tokens in language models (LMs). Thus, “token” and “time slot” will be used interchangeably in the sequel. The vocabulary embedding matrix is denoted as $\bm{E}\in\mathbb{R}^{V\times D}$, where $D$ is the dimension of the time slot embeddings.
We decompose $y_{i}$ into a sequence of tokens $\bm{s_{i}}=\{s_{i}^{1},...,s_{i}^{T_{i}}\}$, where $s_{i}^{t}\in\mathcal{V}$ and $T_{i}$ denotes the length of the sequence. This process, referred to as label encoding, is described in detail in Sec. 3.4. On the other hand, we design a label decoding function $\text{g}(\cdot)$  [^2] Here, $\text{g}(\cdot)$ functions as a lookup table that maps tokens to real-valued vocabulary entries, e.g., $\text{g}(``\text{30s}")=30$. that reconstructs the original watch time $y_{i}$ from $\bm{s_{i}}$, i.e., $y_{i}=\text{g}(\bm{s_{i}})=\sum_{t=1}^{T_{i}}\text{g}(s_{i}^{t})\in\mathbb{R}$. Our goal is to train a sequence generation model, given user and video characteristics $(\bm{u_{i}},\bm{v_{i}})$, which generates the corresponding sequence of watch time slots $\bm{\hat{s}_{i}}=\{\hat{s}_{i}^{1},\hat{s}_{i}^{2},...,\hat{s}_{i}^{T_{i}}\}$, and in turn, from which the predicted watch time $\hat{y_{i}}=\text{g}(\bm{\hat{s}_{i}})=\sum_{t=1}^{T_{i}}\text{g}(\hat{s}_{i}^{t})$ approximates the actual watch time $y_{i}$.
@@ 3.2. The Generative Regression (GR) Model
As shown in Fig. 2, GR adopts a Transformer-based encoder-decoder architecture. The encoder extracts user and video features, while the decoder predicts the watch time in an autoregressive manner.
@@ 3.2.1. Encoder
Unlike traditional sequence-to-sequence tasks or user behavior modeling in recommendation systems, watch time prediction does not inherently depend on the order of user history interacted items. To ensure model generality and simplicity, we follow previous works (Lin et al., 2023; Sun et al., 2024) and employ a feedforward network (FFN) as an encoder. Note that this encoder can be replaced with any sophisticated model architecture. Formally, the encoder extracts user and video features to produce a fixed-length hidden feature $\bm{h_{i}}\in\mathbb{R}^{1\times D}$ that will be fed to the decoder as follows:
@@ 3.2.2. Decoder
where $\theta$ is the model parameter and $\bm{\hat{s}_{i}^{<t}}$ represents the tokens generated before. Utilizing the chain rule, the overall probability of generating the sequence can be expressed as
(3) $P_{\theta}(\bm{s_{i}}\mid\bm{h_{i}})=P_{\theta}(s_{i}^{1},...,s_{i}^{T_{i}}\mid\bm{h_{i}})=\prod_{t=1}^{T}P_{\theta}(s_{i}^{t}\mid\bm{h_{i}},\bm{{s}_{i}^{<t}})$
Three key issues remain to be addressed: (1) how to construct an effective vocabulary, (2) how to encode $y_{i}$ into a sequence $\bm{s_{i}}$, and (3) how to optimize the model. These issues are detailed in the following sections.
@@ 3.3. Vocabulary Construction
  • Completeness: The vocabulary $\mathcal{V}$ must be able to represent all watch time values $\{y_{i}\}_{i=1}^{N}$ using a finite number of tokens almost without loss.Also, each token must be unique.
@@
One intuitive strategy is to select watch time values from the dataset as tokens based on several fixed percentiles, yet failing to meet the completeness principle. An alternative is to select watch time values as tokens based on one fixed percentile, then subtract the token values from all watch time values that exceed them, repeating this process until the residuals become negligible, which fails to meet the balance principle. Due to the space limit, details of this strategy are provided in the supplementary materials.
@@
1:Dataset labels $\bm{Y}=\{y_{j}\}_{j=1}^{N}$, initially empty Vocabulary $\mathcal{V}=\{\}$, start percentile $q_{\text{start}}$, end percentile $q_{\text{end}}$, decay rate $\alpha$, minimal restoration error $\epsilon$.
2:Sort $\bm{Y}$ in descending order to obtain $\hat{\bm{Y}}=\{\hat{y}_{j}\}_{j=1}^{N}$.
@@
5:    Compute the $q$-percentile $o_{i}$ of $\hat{\bm{Y}}$.
6:    if $o_{i}=0$ then $\triangleright$ Terminate if the percentile value is zero
@@
9:    Generate a new token $v_{i}$ which satisfy $o_{i}=g(v_{i})$ and insert $v_{i}$ into vocabulary $\mathcal{V}$.
@@
$\hat{y}_{j}=\begin{cases}\hat{y}_{j},&\text{if }\hat{y}_{j}<o_{i},\\ \hat{y}_{j}-o_{i},&\text{otherwise}\end{cases}$
@@
To address both principles simultaneously, we propose a data-driven vocabulary construction algorithm using dynamic quantile adjustment (Algorithm. 1). The algorithm initializes with a high starting quantile $q_{start}$ and adaptively reduces it by decay rate $\alpha$ until reaching the terminal quantile $q_{end}$. This strategy expedites the reduction of tail values, rapidly decreasing the variance among updated values, which effectively mitigates the challenges posed by the long-tailed distribution in the dataset, for which we provide detailed experimental validation in Sec. 4.4.
We emphasize that our vocabulary construction and label encoding process, while analogous to linguistic syntax building for sequence generation, does not presume theoretical optimality. The proposed strategy serves as a principled engineering solution, leaving theoretical analysis of optimal tokenization for future work.